OpenStudy (anonymous):

x-3 is a factor of P(x)=x³-7x²+15x-9 a. true b. false

OpenStudy (cwrw238):

a

OpenStudy (cwrw238):

as f(3) = 0

jimthompson5910 (jim_thompson5910):

In general, if x-r is a factor of P(x), then P(r) = 0 So simply plug in x = 3 (because here r = 3) and evaluate \[\Large P(x) = x^3-7x^2+15x-9\] \[\Large P(3) = (3)^3-7(3)^2+15(3)-9\] \[\Large P(3) = 27-7(9)+15(3)-9\] \[\Large P(3) = 27-63+45-9\] \[\Large P(3) = 0\] Because P(3) = 0, this means that x-3 is indeed a factor of P(x)